# Premerger phenomena in neutron-star binary coalescences

### v2

@@ IV.3.1 Tidal spinup
In addition to the direct energy deposits from the orbit, tidal interactions torque each of the neutron stars. From a resonant interaction alone, Lai 234 shows that a net angular momentum transfer of
@@
can be expected. An upper bound to the bulk spinup of a given star can then be obtained by assuming that $\Delta J_{\alpha,\rm res}$ contributes only to the uniform rotation of the star, so that $\Delta\Omega_{\alpha,s}\sim\Delta J_{\alpha,\rm res}/I_{0}$. For $f$-modes where the tidal overlap reads $Q\gtrsim 0.1$, the above estimate indicates a non-negligible spinup of order unity may occur, albeit only in the very final moments of merger [see also equation 6.2.6 in 233]. For a $1.4M_{\odot}$ pair of stars, the general result is [e.g. 383]
@@
For $g$- and other non $f$-modes, the effect is likely sub-leading though could achieve $\Delta\Omega_{\alpha,s}/\Omega_{s}\gtrsim 0.01$ for favourable overlaps. The long-lived tidal-grinding provided by the non-resonant $f$-mode could also be significant. Either way, such effects complicate the dynamics of the system considerably, as one must, in general, account for a time-dependent spin factor in modelling the tidal dynamics. Such modelling has not been attempted aside from the work of Lai 234 for neutron stars, and in the different contexts of binaries involving a white dwarf by Racine et al. 324, Fuller and Lai 149.
@@ IV.4 Magnetic effects
Magnetic fields control all of the electromagnetic phenomena associated with neutron stars, including those occurring premerger (see Secs. II.3 and III.2). Dynamical magnetic fields can also generate electric fields via induction which work towards accelerating the binary inspiral [197].
As described by Lai 235 and others prior [beginning with 163], suppose that the magnetic field of the primary greatly exceeds of the companion, so that the latter may be treated effectively as a perfect conductor with vanishing dipole moment. The orbital motion of this companion through the rotating magnetosphere of the primary generates an electromotive force, driving a “direct current” between the constituents, with the closed magnetic field lines playing the role of wires in the circuit. This is the so-called “unipolar inductor” model, which will be covered in some more detail in Sec. VI.2. The exact way in which EM energy is siphoned out of the orbit depends on the total resistance of the circuit, $\mathcal{R}$, which includes contributions from the primary, companion, and the (intertwined) magnetosphere. Lai 235 argues that the maximum energy dissipation rate reads
@@
for electromotive torque $T_{\rm mag}$, primary (equatorial) dipole moment $\mu_{d}$, and twist parameter $\zeta_{\phi}$. This latter parameter relates to the induced toroidal field due to kinetics, as is familiar from studies of low mass X-ray binaries [e.g. 162], though its value is highly uncertain: it depends on poorly-understood boundary-layer physics.
In fact, Piro 312 suggested implicitly (i.e. in a different notation) that the twist could reach values much larger than unity, $\zeta_{\phi}\gg 1$, which could significantly accelerate the inspiral. Furthermore, the torque $T_{\rm mag}$ will generally spin up the primary when $\Omega_{\rm orb}>\Omega_{s}$; Piro 312 argued this could lead to near synchronization, and thus rapidly rotating stars prior to merger if the dipole moment is sufficiently large (cf. Sec. II.2). On the other hand, if the magnetic pressure exerted by the toroidal field is too large, the flux tubes defining the circuit will be broken by unstable kink modes. This is the essence of the argument put forward by Lai 235 regarding the limiting value of the twist ($\zeta_{\phi}\sim 1$), and hence a relatively weak electromagnetic contribution (47). Still, quasi-periodic circuits may be established if after such a flux tube breaks, reconnection between the (inflated) field lines re-links the two stars so that the cycle may repeat. Understanding the dynamics of such a system requires 3D simulations which have not yet been conducted; though see Crinquand et al. 100, who find that $\zeta_{\phi}$ could greatly exceed unity in cases where the stellar spins are anti-aligned with respect to each other.
@@
together with $\Omega_{\rm orb}^{2}\approx GM(1+q)/a^{3}$ we can evaluate expression (49) easily. One set of results for various values of $\zeta_{\phi}$ are shown in Figure 11, where we postulate a magnetar-level field at the equator of $\sim 10^{14}$ G, radius $12$ km, moment of inertia $I_{0}=10^{45}\text{ g cm}^{2}$, and two $1.4M_{\odot}$ stars ($q=1$). The evolution is tracked starting from a separation of $200$ km ($\Omega_{\rm orb}\approx 216$ Hz). The magnetic star is practically static initially with $\Omega_{s,0}=0.1$ Hz, though half a second before merger ($a\approx 9R$) can reach non-negligible spins if the twist parameter is large ($\zeta_{\phi}\gtrsim 1$). For the case $\zeta_{\phi}=10^{2}$ — still significantly less than that estimated by Piro 312 for some systems — the spin frequency could reach $\nu\gtrsim 1$ Hz when $a(t)\sim 9R$. Larger field strengths amplify the effect further like $\sim B^{2}$. If some magnetars with larger fields take part in mergers, as hinted at by some precursor observations (see Sec. VI), spin-up could be dynamically impactful. These spinups will combine with that of the tidal torques discussed in Sec. IV.3.1.
@@ IV.4.1 Internal fields?
Magnetic effects can also be associated with that of the interior, rather than the exterior. A strong tension can either reduce or increase the tidal deformability of a star depending on the nature of the magnetic geometry: poloidal (toroidal) fields tend to induce an oblate (prolate) distortion. This was studied by Giacomazzo et al. 157 and more recently by Zhu et al. 450 using complementary techniques, though the results are rather pessimistic: relative shifts in the deformabilities $\Lambda$ (see below equation 39) amount to $\lesssim 0.1\%$ for $B\lesssim 10^{15}$ G for neutron-star EOS. For quark-star EOS, the effects are greater. In a superconducting core, the result may also be somewhat larger if $H_{1}\sim 10^{16}$ G [e.g. 239, and equation 10], which leads to a factor $\gtrsim 10$ increase in the effective magnetic tension and hence a percent-level shift for magnetars. Tidal deformabilities for stars with superconducting components have not been calculated in the literature. Some additional effects could arise if the internal magnetic fields are large enough to distort the stellar EOS, though these must be typically larger than even $H_{c1}$ [see, e.g. 118, 379]. A recent discussion on mathematical elements of GRMHD equilibria can be found in Suvorov and Glampedakis 380.
@@ IV.5 Thermal and viscous effects
As quantified by equation (25), tidal energy is siphoned into QNMs during inspiral. This energy is essentially in the form of the potential and kinetic energy associated with the master eigenfunction $\boldsymbol{\xi}$. As detailed by Goldreich and Reisenegger 164 and Lai 235, three ways in which these modes can be damped is through GW radiation-reaction, chemical relaxation through neutrino emission, and viscous dissipation. The first of these is largely irrelevant (though note the damping time of $f$-modes can be of the order $\sim 100$ ms, which could theoretically be fast enough if a very early resonance occurs because of rapid rotation; see Sec. IV.3).
Viscous heating is likely to play a significant role. As described in Kuan et al. 232 and elsewhere, the stellar stratification is generally a function of time because of this heating, implying that the $g$-mode spectrum itself is time-dependent [this gets even more complicated if accounting for nuclear reactions; 176]. In particular, Lai 233 has shown that the non-resonant $\ell=2=m$ $f$-mode in an aligned (or non-rotating) binary will increase the star’s temperature by (see equation 8.30 therein)
@@
It has been suggested that $T$ could reach values considerably larger than (51). Arras and Weinberg 39 — who also provide an overview of the tidal heating literature in their Introduction — find that chemical heating, driven by the induced fluid motions during inspiral, could raise the temperature to $\sim 2\times 10^{8}$ K if the system has access to the direct Urca mechanism. This is because density perturbations excited via tides force the system out of beta equilibrium, instigating chemical heating. If the core contains “strange” matter such that nonleptonic weak interactions raise the bulk viscosity by a couple orders of magnitude (e.g. hyperonic matter), tidal heating could even raise the temperature to $\lesssim 10^{10}$ K [156]. Such enormous temperatures would have a significant impact on the $g$-spectrum of the system leading to dephasing, aside from that which is associated with the heating directly. The latter effect would likely exceed $\sim 0.1$ rad and thus be detectable with the LVK network [156]; see also Sec. IV.5.1. In the limit of totally efficient tidal dissipation the core temperature would easily reach $\sim 10^{10}$ K before merger, as calculated by Meszaros and Rees 270.
@@
Overall, thermal elements in mergers are rather subtle and not fully understood beyond the above. For example, Hammond et al. 176 point out that out-of-equilibrium elements in a merger lead to a softening of the EOS in some density regions, and to composition changes that affect processes that rely on deviation from equilibrium, such as bulk viscosity, both in terms of the magnitude and the equilibriation timescales inherent to the relevant set of reaction rates. Whether the estimates described above may apply in a realistic merger environment, including the possibility of constraining them, is not obvious. Significant tidal heating of the stars can also induce mass loss via winds before merger [270], which could lead to electromagnetic, premerger emissions if strong shocks develop in the winds. Such debris will also contribute to the dynamical and postmerger ejecta polluting the eventual crash site, which could inhibit jet formation and affect the breakout timescale (see Sec. VI.1). If the stars reach high-enough premerger temperatures, the bulk viscous timescales calculated in [19] could be utilised.
@@ IV.6 Residual eccentricity
Aside from canonical formation channels involving a symbiotic binary (see Sec. II.2), mergers might occur in globular clusters where captures are not irregular. Although the recent Monte Carlo simulations of Ye et al. 439 estimate a globular-cluster merger rate of only $\sim 0.02\text{ Gpc}^{-3}\text{ yr}^{-1}$, such systems may retain a significant eccentricity as the usual circularisation arguments made in Sec. IV following the Peters 309 formulae do not apply. Tidal phenomena in eccentric inspirals acquire additional complexity because resonances can be triggered at different orbital phases (likely near periastron), and thus generally will happen earlier; multiple resonances can also be triggered each passage.
It has been shown that for eccentric binaries there can be chaotic growth of the modes, with the kinetic energy stored in a mode the rivalling that of the orbital binding energy, $U\sim GMm/a$, after many cycles [258, 415]. Given that this energy greatly exceeds that which is typically expected to be stored in resonant modes, $E_{\rm kin,max}\sim 10^{45}$erg, it would be worthwhile to revisit such studies but for neutron stars specifically [89, 438, 423]. One such study was carried out by Vick and Lai 416, finding appreciable effects depending on a complicated relationship between the eccentricity and how much energy is siphoned into the $f$-mode; see also Takátsy et al. 394.
@@ IV.7 Remarks on merger simulations and future challenges
Here we provide a brief discussion on merger simulation results with respect to mode excitations; we direct the reader interested in GRMHD details more generally to Baiotti and Rezzolla 44, Kiuchi 211. Although the mode evolution is not as transparent as in Figs. 810, some fully relativistic simulations may resolve the excitation of the $f$-mode in the last $<100$ ms of inspiral given that the dephasing between the numerical waveform and the analytic model including realistic $f$-mode effects is small [366, 128, 154]. Aside from the $f$-mode, numerically capturing the excitations of other (polar) modes that could experience a resonance (viz. $g$-, $i$-, shear-, and ocean-modes) are currently out of reach due to a number of technical and theoretical challenges. This is important since $g$-mode dephasings in particular may not be totally negligible [discussed recently by 189, for instance] and represents a future challenge for the numerical-relativity community. Failing to account for such dephasings, if large, could lead to spurious inferences about neutron-star structure.
@@
Spatial resolution is not the only technical requirement to see the excitation of lower frequency modes. In such cases, one would need to start the simulation at a time when the orbital frequency is less than $\sim 50$ Hz to include their resonance windows. This corresponds to many seconds prior to merger (cf. Fig. 8). This is borderline impossible owing to both computational demand and numerical stability considerations when coupled in with the above resolution demand, at least in a full numerical-relativity sense where the spacetime is evolved self-consistently with (magneto-)hydrodynamics using the full Einstein equations. For higher quantum numbers ($g_{n\geq 2}$, …) even greater resolution is required, and the problem becomes more severe still. On top of this, $g$- and other modes’ tidal response tend to be rather shallow compared to the $f$-mode [see 233, 217, 353, 302, and others], and thus numerical dissipation can easily bias the results, i.e. convergence is expected to be much harder to obtain than the $f$-mode case. This is in addition to the usual separation problem of the real and imaginary components of $g$-modes [e.g. 230].
@@
From a theoretical perspective, most (though not all) of the long-term inspiral simulations adopt a cold EOS through a piecewise-polytropic approximation which does not include the crust structure and cannot support the thermal and compositional stratification gradients that exist in a real star. Therefore, all of the microphysically-dependent modes may not even exist in the first place within the computational setups employed. It is thus even difficult to estimate how many CPU hours and further code innovations would be needed to accurately simulate the evolution of these modes and claim it was either resolved or not resolved [cf. 176]. The highly-dynamical nature of the spacetime and fluid makes it also not obvious whether “modes” even exist in the usual sense of the word, as these are inherently linear while the Einstein equations are inherently nonlinear [see, e.g. 152, 360, for discussions].

### v3

@@ IV.3.1 Tidal spinup
In addition to the direct energy deposits from the orbit, tidal interactions torque each of the neutron stars. From a resonant interaction alone, [243] shows that a net angular momentum transfer of
@@
can be expected. An upper bound to the bulk spinup of a given star can then be obtained by assuming that $\Delta J_{\alpha,\rm res}$ contributes only to the uniform rotation of the star, so that $\Delta\Omega_{\alpha,s}\sim\Delta J_{\alpha,\rm res}/I_{0}$. For $f$-modes where the tidal overlap reads $Q\gtrsim 0.1$, the above estimate indicates a non-negligible spinup of order unity may occur, albeit only in the very final moments of merger [see also equation (6.2.6) in [242]]. For a $1.4M_{\odot}$ pair of stars, the general result is [e.g. 397]
@@
For $g$- and other non $f$-modes, the effect is likely sub-leading though could achieve $\Delta\Omega_{\alpha,s}/\Omega_{s}\gtrsim 0.01$ for favourable overlaps. The long-lived tidal-grinding provided by the non-resonant $f$-mode could also be significant. Either way, such effects complicate the dynamics of the system considerably, as one must, in general, account for a time-dependent spin factor in modelling the tidal dynamics. Such modelling has not been attempted aside from the work of [243] for neutron stars, and in the different contexts of binaries involving a white dwarf by [337, 152].
@@ IV.4 Magnetic effects
Magnetic fields control all of the electromagnetic phenomena associated with neutron stars, including those occurring premerger (see Secs. II.3 and III.2). Dynamical magnetic fields can also generate electric fields via induction which work towards accelerating the binary inspiral.
As described by Lai 244 and others prior [beginning with 168], suppose that the magnetic field of the primary greatly exceeds of the companion, so that the latter may be treated effectively as a perfect conductor with vanishing dipole moment. The orbital motion of this companion through the rotating magnetosphere of the primary generates an electromotive force, driving a “direct current” between the constituents, with the closed magnetic field lines playing the role of wires in the circuit. This is the so-called “unipolar inductor” model, which will be covered in some more detail in Sec. VI.2. The exact way in which EM energy is siphoned out of the orbit depends on the total resistance of the circuit, $\mathcal{R}$, which includes contributions from the primary, companion, and the (intertwined) magnetosphere. Lai 244 argues that the maximum energy dissipation rate reads
@@
for electromotive torque $T_{\rm mag}$, primary (equatorial) dipole moment $\mu_{d}$, and twist parameter $\zeta_{\phi}$. This latter parameter relates to the induced toroidal field due to kinetics, as is familiar from studies of low mass X-ray binaries [e.g. 167], though its value is highly uncertain: it depends on poorly-understood boundary-layer physics.
In fact, Piro 325 suggested implicitly (i.e. in a different notation) that the twist could reach values much larger than unity, $\zeta_{\phi}\gg 1$, which could significantly accelerate the inspiral. Furthermore, the torque $T_{\rm mag}$ will generally spin up the primary when $\Omega_{\rm orb}>\Omega_{s}$; Piro 325 argued this could lead to near synchronization, and thus rapidly rotating stars prior to merger if the dipole moment is sufficiently large (cf. Sec. II.2). On the other hand, if the magnetic pressure exerted by the toroidal field is too large, the flux tubes defining the circuit will be broken by unstable kink modes. This is the essence of the argument put forward by Lai 244 regarding the limiting value of the twist ($\zeta_{\phi}\sim 1$), and hence a relatively weak electromagnetic contribution (47). Still, quasi-periodic circuits may be established if after such a flux tube breaks, reconnection between the (inflated) field lines re-links the two stars so that the cycle may repeat. Understanding the dynamics of such a system requires 3D simulations which have not yet been conducted; though see Crinquand et al. 99, who argue via numerical simulations that $\zeta_{\phi}$ could greatly exceed unity in cases where the stellar spins are anti-aligned with respect to each other.
@@
together with $\Omega_{\rm orb}^{2}\approx GM(1+q)/a^{3}$ we can evaluate expression (49) easily. One set of results for various values of $\zeta_{\phi}$ are shown in Figure 11, where we postulate a magnetar-level field at the equator of $\sim 10^{14}$ G, radius $12$ km, moment of inertia $I_{0}=10^{45}\text{ g cm}^{2}$, and two $1.4M_{\odot}$ stars ($q=1$). The evolution is tracked starting from a separation of $200$ km ($\Omega_{\rm orb}\approx 216$ Hz). The magnetic star is practically static initially with $\Omega_{s,0}=0.1$ Hz, though half a second before merger ($a\approx 9R$) can reach non-negligible spins if the twist parameter is large ($\zeta_{\phi}\gtrsim 1$). For the case $\zeta_{\phi}=10^{2}$ — still significantly less than that estimated by Piro 325 for some systems — the spin frequency could reach $\nu\gtrsim 1$ Hz when $a(t)\sim 9R$. Larger field strengths amplify the effect further like $\sim B^{2}$. If some magnetars with larger fields take part in mergers, as hinted at by some precursor observations (see Sec. VI), spin-up could be dynamically impactful. These spinups will combine with that of the tidal torques discussed in Sec. IV.3.1.
@@ IV.4.1 Internal fields?
Magnetic effects can also be associated with that of the interior, rather than the exterior. A strong tension can either reduce or increase the tidal deformability of a star depending on the nature of the magnetic geometry: poloidal (toroidal) fields tend to induce an oblate (prolate) distortion. This was studied by [161] and more recently by [465] using complementary techniques, though the results are rather pessimistic: relative shifts in the deformabilities $\Lambda$ [see below equation (39)] amount to $\lesssim 0.1\%$ for $B\lesssim 10^{15}$ G for neutron-star EOS. For quark-star EOS, the effects are greater. In a superconducting core, the result may also be somewhat larger if $H_{1}\sim 10^{16}$ G [e.g. [248] and equation (10)], which leads to a factor $\gtrsim 10$ increase in the effective magnetic tension and hence a percent-level shift for magnetars. Tidal deformabilities for stars with superconducting components have not been calculated in the literature. Some additional effects could arise if the internal magnetic fields are large enough to distort the stellar EOS, though these must be typically larger than even $H_{c1}$ [see, e.g. 118, 393]. A recent discussion on mathematical elements of GRMHD equilibria can be found in Suvorov and Glampedakis [394].
@@ IV.5 Thermal and viscous effects
As quantified by equation (25), tidal energy is siphoned into QNMs during inspiral. This energy is essentially in the form of the potential and kinetic energy associated with the master eigenfunction $\boldsymbol{\xi}$. As detailed by Lai 244 and others, three ways in which these modes can be damped is through GW radiation-reaction, chemical relaxation through neutrino emission, and viscous dissipation. The first of these is largely irrelevant (though note the damping time of $f$-modes can be of the order $\sim 100$ ms, which could theoretically be fast enough if a very early resonance occurs because of rapid rotation; see Sec. IV.3).
Viscous heating is likely to play a significant role. As described in Kuan et al. 241 and elsewhere, the stellar stratification is generally a function of time because of this heating, implying that the $g$-mode spectrum itself is time-dependent [this gets even more complicated if accounting for nuclear reactions; 182]. In particular, Lai 242 has shown that the non-resonant $\ell=2=m$ $f$-mode in an aligned (or non-rotating) binary will increase the star’s temperature by (see equation 8.30 therein)
@@
It has been suggested that $T$ could reach values considerably larger than (51). Arras & Weinberg [39] — who also provide an overview of the tidal heating literature in their Introduction — find that chemical heating, driven by the induced fluid motions during inspiral, could raise the temperature to $\sim 2\times 10^{8}$ K if the system has access to the direct Urca mechanism. This is because density perturbations excited via tides force the system out of beta equilibrium, instigating chemical heating. If the core contains “strange” matter such that nonleptonic weak interactions raise the bulk viscosity by a couple orders of magnitude (e.g. hyperonic matter), tidal heating could even raise the temperature to $\lesssim 10^{10}$ K [160]. Such enormous temperatures would have a significant impact on the $g$-spectrum of the system leading to dephasing, aside from that which is associated with the heating directly. The latter effect would likely exceed $\sim 0.1$ rad and thus be detectable with the LVK network [160]; see also Sec. IV.5.1. In the limit of totally efficient tidal dissipation the core temperature would easily reach $\sim 10^{10}$ K before merger, as calculated by Meszaros and Rees [281].
If the system were to reach even a fraction of such values, the effective Love number should decrease because extra thermal pressures cause the outer boundary ($p=0$) to extend. Using various EOS, Kanakis-Pegios, Koliogiannis, & Moustakidis [211] calculated the quadrupolar Love numbers at fixed temperatures finding that for $T\sim 1$ MeV ($\sim 10^{10}$ K) there is a decrease in $k_{2}$ of order $\approx 20\%$ with a $\approx 70\%$ drop for $T\sim 5$ MeV; the tidal deformability, however, is largely insensitive to the temperature since the radius increases to compensate the reduced $k_{2}$ for the EOS considered therein. High temperatures would also impact on the elastic maximum of the crust (Sec. III.2).
@@
Overall, thermal elements in mergers are rather subtle and not fully understood beyond the above. For example, it is pointed out in [182] that out-of-equilibrium elements in a merger lead to a softening of the EOS in some density regions, and to composition changes that affect processes that rely on deviation from equilibrium, such as bulk viscosity, both in terms of the magnitude and the equilibriation timescales inherent to the relevant set of reaction rates. Whether the estimates described above may apply in a realistic merger environment, including the possibility of constraining them, is not obvious. Significant tidal heating of the stars can also induce mass loss via winds before merger [281], which could lead to electromagnetic, premerger emissions if strong shocks develop in the winds. Such debris will also contribute to the dynamical and postmerger ejecta polluting the eventual crash site, which could inhibit jet formation and affect the breakout timescale (see Sec. VI.1). If the stars reach high-enough premerger temperatures, the bulk viscous timescales calculated in [19] could be utilised.
@@ IV.6 Residual eccentricity
Aside from canonical formation channels involving a symbiotic binary (see Sec. II.2), mergers might occur in globular clusters where captures are not irregular. Although the recent Monte Carlo simulations of Ye et al. 453 estimate a globular-cluster merger rate of only $\sim 0.02\text{ Gpc}^{-3}\text{ yr}^{-1}$, such systems may retain a significant eccentricity as the usual circularisation arguments made in Sec. IV following the Peters 322 formulae do not apply. Tidal phenomena in eccentric inspirals acquire additional complexity because resonances can be triggered at different orbital phases (likely near periastron), and thus generally will happen earlier; multiple resonances can also be triggered each passage.
It has been shown that for eccentric binaries there can be chaotic growth of the modes, with the kinetic energy stored in a mode the rivalling that of the orbital binding energy, $U\sim GMm/a$, after many cycles [267, 429]. Given that this energy greatly exceeds that which is typically expected to be stored in resonant modes, $E_{\rm kin,max}\sim 10^{45}$erg, it would be worthwhile to revisit such studies but for neutron stars specifically [88, 452, 437]. One such study was carried out by Vick and Lai 430, finding appreciable effects depending on a complicated relationship between the eccentricity and how much energy is siphoned into the $f$-mode; see also Takátsy et al. 408.
@@ IV.7 Remarks on merger simulations and future challenges
Here we provide a brief discussion on merger simulation results with respect to mode excitations; we direct the reader interested in GRMHD details more generally to [44, 218]. Although the mode evolution is not as transparent as in Figs. 810, some fully relativistic simulations may resolve the excitation of the $f$-mode in the last $<100$ ms of inspiral given that the dephasing between the numerical waveform and the analytic model including realistic $f$-mode effects is small [380, 129, 157]. Aside from the $f$-mode, numerically capturing the excitations of other (polar) modes that could experience a resonance (viz. $g$-, $i$-, shear-, and ocean-modes) are currently out of reach due to a number of technical and theoretical challenges. This is important since $g$-mode dephasings in particular may not be totally negligible (discussed recently by [195] for instance) and represents a future challenge for the numerical-relativity community. Failing to account for such dephasings, if large, could lead to spurious inferences about neutron-star structure.
@@
Spatial resolution is not the only technical requirement to see the excitation of lower frequency modes. In such cases, one would need to start the simulation at a time when the orbital frequency is less than $\sim 50$ Hz to include their resonance windows. This corresponds to many seconds prior to merger (cf. Fig. 8). This is borderline impossible owing to both computational demand and numerical stability considerations when coupled in with the above resolution demand, at least in a full numerical-relativity sense where the spacetime is evolved self-consistently with (magneto-)hydrodynamics using the full Einstein equations. For higher quantum numbers ($g_{n\geq 2}$, …) even greater resolution is required, and the problem becomes more severe still. On top of this, $g$- and other modes’ tidal response tend to be rather shallow compared to the $f$-mode [see 242, 225, 367, 315, and others], and thus numerical dissipation can easily bias the results, i.e. convergence is expected to be much harder to obtain than the $f$-mode case. This is in addition to the usual separation problem of the real and imaginary components of $g$-modes [e.g. 239].
@@
From a theoretical perspective, most (though not all) of the long-term inspiral simulations adopt a cold EOS through a piecewise-polytropic approximation which does not include the crust structure and cannot support the thermal and compositional stratification gradients that exist in a real star. Therefore, all of the microphysically-dependent modes may not even exist in the first place within the computational setups employed. It is thus even difficult to estimate how many CPU hours and further code innovations would be needed to accurately simulate the evolution of these modes and claim it was either resolved or not resolved (cf. [182]). The highly-dynamical nature of the spacetime and fluid makes it also not obvious whether “modes” even exist in the usual sense of the word, as these are inherently linear while the Einstein equations are inherently nonlinear [see, e.g. 155, 374, for discussions].